A charge particle ‘q’ is shot towards another charged particle ‘Q’ which is fixed, with a speed ‘v’ . It approaches ‘q’ upto a closest distance r and then returns. If q were given a speed of ‘2v’ the closest distances of approach would be

Grade:6

4 Answers

Navjyot Kalra
9 years ago
(d) ½ mv2 = kQq/r ? ½ m (2v)2 = kqQ/r’ ? r’ = r/4
shubham
6 years ago
as mentioned by navjyot 
all the kinetic energy is being completely converted into potential energy of system of two(both) particles
 
swashna
5 years ago
as per the kinetic energy formula.statement Navjyot is correct.
her the kinetic energy is totally converted into potential energy
ankit singh
3 years ago
= kqQ/r’ ? r’ =r/4(by the eneregy conservation equation).......................................2 = kQq/r ? ½ m (2v)2½ mv